OpenStudy (anonymous):

Evaluate: −6 − (−4) −10 −2 24 2

OpenStudy (anonymous):

-6+4= -2 so -2 is the answer

OpenStudy (anonymous):

−6 − (−4) = -6+4 now do the last step

OpenStudy (anonymous):

i cant

OpenStudy (anonymous):

why???

OpenStudy (anonymous):

cause i dont know how to do this pellet.

OpenStudy (anonymous):

haha in which grade you are???

OpenStudy (anonymous):

7

OpenStudy (anonymous):

so you must understand this. its very easy

OpenStudy (anonymous):

i figured it out